OpenStudy (anonymous):

calculate the equation of the line that is parallel to the line y=4x-6 and passes through the point (5,8)

OpenStudy (anonymous):

hint-parallel lines have same slopes.

OpenStudy (anonymous):

what is the slope of y=4x-6?

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

yes., now the equation of the required line is y=4x+c. do you agree?

OpenStudy (texaschic101):

It will have the same slope because it is a parallel line. So we have a slope of 4. Now we will use this formula : y - y1 = m(x - x1) m = 4 y1 = 8 x1 = 5 now lets sub y - 8 = 4(x - 5) y - 8 = 4x - 20 y = 4x - 20 + 8 y = 4x - 12 <-- your equation